# GraspLDM: Generative 6-DoF Grasp Synthesis using Latent Diffusion Models

### v1

@@ Abstract
Vision-based grasping of unknown objects in unstructured environments is a key challenge for autonomous robotic manipulation. A practical grasp synthesis system is required to generate a diverse set of 6-DoF grasps from which a task-relevant grasp can be executed. Although generative models are suitable for learning such complex data distributions, existing models have limitations in grasp quality, long training times, and a lack of flexibility for task-specific generation. In this work, we present GraspLDM- a modular generative framework for 6-DoF grasp synthesis that uses diffusion models as priors in the latent space of a VAE. GraspLDM learns a generative model of object-centric $SE(3)$ grasp poses conditioned on point clouds. GraspLDM’s architecture enables us to train task-specific models efficiently by only re-training a small de-noising network in the low-dimensional latent space, as opposed to existing models that need expensive re-training. Our framework provides robust and scalable models on both full and single-view point clouds. GraspLDM models trained with simulation data transfer well to the real world and provide an 80% success rate for 80 grasp attempts of diverse test objects, improving over existing generative models. We make our implementation available at https://github.com/kuldeepbrd1/graspldm.
@@ I Introduction
Robotic grasping of unknown objects from visual observations is an integral skill for autonomous manipulation systems in the real world. Grasping is the fundamental manipulation task of generating restraining contacts on an object. An essential part of this task is grasp synthesis, which involves reasoning about a set of grasps that are possible around an object. From this set of possible grasps, a subjectively good grasp can then be chosen based on grasp quality, task context, and kinematic feasibility. Vision-based grasp synthesis is a challenging problem for three main reasons. First, there are infinite possible grasps on any object and the distribution of good grasps can be complex, with multiple modes and discontinuities. Second, the grasp generation process has to rely on unstructured visual observations from the sensors. Finally, it needs to generalize to an arbitrary set of unknown objects that a robot may encounter in the real world. To alleviate some of these issues, recent works [1] have proposed to use learning-based approaches. While generating 6-Degrees of Freedom (DoF)poses remains difficult [2], a simplification of this problem to tabletop objects, leading to a 4-DoF grasping space, has shown promising results.
Figure 1: GraspLDM uses a de-noising diffusion model in the latent space of a VAE to improve grasp generation performance. It also enables injection of task-conditional guidance in a modular manner.
@@
In this work, we focus on generating object-centric 6-DoF grasps using point clouds. We assume a parallel jaw gripper and parameterize a grasp by its $SE(3)$ pose. The objective is to learn a representation of object-centric 6-DoF grasps that can generalize to unknown objects. Learning this representation is difficult, as the model must reason about an arbitrary set of grasp poses for each object in the $SE(3)$ space. Deep generative models are suitable for this problem, as they are designed to learn complex data-generating distributions efficiently. 6DoF-Graspnet [3], a primary generative grasp synthesis model, learns the distribution of object-centric grasps in a continuous latent space of a conditional Variational Autoencoder (VAE). However, the VAE exhibits poor sampling quality, which can be attributed to the prior gap problem in VAE s when the encoding distribution does not match the prior very well [4]. On the other hand, diffusion models [5, 6] are a class of generative models that can overcome these problems. In particular, the latent diffusion [7] formulation provides an efficient and flexible generation architecture by using a diffusion model in the latent space of a VAE.
@@
In this work, we propose GraspLDM - a learning-based generative modeling approach for 6-DoF grasp synthesis. GraspLDM learns the distribution of successful grasps on object point clouds using latent diffusion similar to [8], that transfers to the real world (Fig. 0). In GraspLDM, a diffusion model is trained efficiently inside the low-dimensional latent space of a VAE. This diffusion model acts as a prior and bridges the gap between the VAE prior and the posterior distribution fit during the training. Using diffusion in the latent space improves the quality of sampled grasps and avoids the representational complexity of operating directly on pairs of point clouds and grasp pose. It also enables two aspects of model flexibility that are not available in grasp synthesis models based on VAE [3] and diffusion [9]. First, task-specific conditional guidance can be provided solely in the latent space without re-training the VAE encoder and decoder. Second, multiple diffusion models can be trained efficiently and plugged inside this low-dimensional latent space.
@@ II Related Work
Grasp synthesis was already studied by [10] using analytical methods, which rely on the knowledge of object models and properties that are rarely available in the real world. Today, grasp synthesis is framed as the problem of finding stable grasp poses from sensor data, subject to a gripper. Data-driven and learning-based methods have made notable progress, especially on 4-DoF planar grasp synthesis using parallel and vacuum grippers. In contrast, generalizable 6-DoF grasp synthesis is an active research topic [2] with two broad classes: discriminative and generative approaches. Discriminative approaches rely on a manual pose sampling strategy and learn to discriminate good grasps from bad grasps using a quality measure [11]. Generative approaches instead directly learn to generate grasp poses from an observation [12, 3] with implicit sampling. Since grasps can be sampled directly from the learned model [3] these tend to be efficient. Some works on 6-DoF grasping also address auxiliary tasks to improve grasp generation, like shape completion and 3D reconstruction [13]. In this work, we present a generative approach that learns a continuous distribution of object-centric grasps conditioned on point clouds in an end-to-end manner.
@@ II-C Denoising Diffusion Models
GraspLDM utilizes a diffusion model based on Denoising Diffusion Probabilistic Models (DDPM)[6]. Score-based Generative Model (SGM[5] and Stochastic Differential Equations (SDE)are equivalent formulations of the diffusion process with notable implementation differences. For grasp pose generation, [9] uses the SGM formulation and learns a scalar field representing the denoising score. While their models effectively refine randomly sampled SE(3) grasp poses to low-cost (good grasp) regions, the analysis only presents learning on full point clouds of a single category (Mugs). Real-world tests on Mugs in [9] use ground truth pose of the object. Further, Langevin-type sampling is slow and the model needs to be retrained for downstream tasks. In contrast, GraspLDM is a more flexible formulation to [9] that operates a diffusion model on a vector field in the continuous latent space of a VAE. Our analyses and validation are also more extensive using full and partial point clouds from a large object category set, complemented with real-world transfer tests. GraspLDM can also use fast samplers like Denoising Diffusion Implicit Model (DDIM)as a drop-in replacement post-hoc with minimal performance loss. Further, GraspLDM can be extended to gripper parameterizations beyond the pose, in contrast to [9]. Diffusion models have also been utilized for dexterous grasp synthesis [16], but rely on the availability of the object model.
@@ III Grasp Latent Diffusion Models
Since we do not have access to the real latent space, computing the likelihood of data over the whole latent space in Eq. 1 is intractable. To overcome this, we use a VAE formulation for learning our generative model. VAE’s encoder compresses the complex point cloud-grasp pair input to a continuous latent space with a prior distribution. However, the approximate posterior of the encoder does not match the prior perfectly, especially if it prioritizes data reconstruction for grasp accuracy. To tackle this, we use a diffusion model in its latent space. Together, they form the latent diffusion model for grasp generation that we call GraspLDM. The diffusion model simultaneously improves sample quality by bridging the prior gap in the latent space and enables the introduction of latent space guidance with task context. The following sections describe the GraspLDM design in detail.
@@ III-A Conditional Variational Autoencoder
In Eq. 2, the first term is the likelihood of the decoder reconstructing the input data. The second term measures the divergence between the approximate posterior distribution $q_{\psi}(\mathbf{z}_{h}|H,\mathbf{x}_{pc})$ of the encoder and the assumed prior, $\mathbf{z}\sim\mathcal{N}(0,1)$. $\lambda$ is the weighting hyperparameter that controls the strength of KL regularization in the latent space. A constant $\lambda=1$ provides the strict ELBO objective. However, this frequently results in KL divergence term ($D_{KL}$) becoming vanishingly small early in the training. When this happens, the conditional decoder becomes auto-regressive and the latent has no effect on the output [4]. To avoid this, we use linear annealing on the $\lambda$ parameter.
@@ III-B Latent Diffusion
In principle, sampling from VAE’s prior and decoding it should be sufficient to generate successful grasps. However, decoding a latent drawn from the prior often results in lower sample quality because of the prior gap problem highlighted earlier. To alleviate this issue, we propose to use a Denoising Diffusion Model (DDM)in the latent space. We justify the choice of latent diffusion structure from two observations. First, a simple DDPM would need to encode a point cloud at each step and use a high-dimensional intermediate representation, increasing computation for each denoising step. Alternatively, separating the point cloud encoder from the diffusion requires separate expensive training of a point cloud auto-encoder. Further, such a model needs to be re-trained for newer task conditionings.
A DDM consists of two processes: a forward diffusion process and a reverse de-noising process. In this work, we use the standard discrete-time DDPM [6], where the forward process is a linear Markov chain of deterministic Gaussian kernels in Eq. 3a. The noise added at each time-step $t\in[1,T]$ is pre-defined by a variance schedule $\beta_{t}$. We use linearly increasing $\beta_{t}$ such that the distribution $q(\mathbf{z}_{T})$ converges to the standard normal distribution at the final forward time step $T$. In the forward diffusion process, a noisy sample at any time step can then be obtained by Eq. 3b.
@@
Intuitively, the forward diffusion process adds noise to input data until the data distribution transitions to a simple prior distribution. Consequently, recovering the data distribution from the prior requires a time reversal of the forward process from $t=T$ to $t=0$. This reverse diffusion process uses a learned score network ($\theta$) which de-noises a sample, in our case the latent $\mathbf{z}_{h}$, from $t$ to $t-1$. For notational simplicity, we use $\mathbf{z}_{t}$ to refer to conditional grasp latent $\mathbf{z}_{h}$ at time $t$. Using the reverse diffusion kernel in Eq. 4, the distribution of the de-noised sample at time-step $t-1$ can be modeled by the mean $\mu_{\theta}(z_{t},t)$ and a known variance $\sigma_{t}^{2}$. More simply, we can re-parameterize the mean and instead learn a network to directly predict the noise $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t)$ to be removed at each step [6]. The de-noising score model $\bm{\epsilon}_{\theta}$ predicts the noise to be removed conditioned on the current time-step, while the parameter weights ($\theta$) are shared across all the time steps. Then, the sampling can be done using Eq. 5, where $\sigma_{t}$ can be simplified to use variance similar to the forward process i.e. $\beta_{t}$. DDMs with this formulation can be trained to maximize an ELBO-like objective. Given the simplified formulation in [6] and adapting it to our latent diffusion architecture, this objective $\mathcal{L}_{D}(\theta)$ can be written as Eq. 6.
@@
((a))
((b))
Figure 3: Grasp generation performance and scaling on full object point clouds ($N=1024$). (a) The mean success rate in simulation of 300 generated grasps poses per object. (b) SE(3) EMD between ground-truth grasp pose distribution and 100 sampled grasp poses (lower is better). SE3-DiF-1C and SE3-DiF-63C are the SE(3) Grasp Diffusion models from [9].
@@ III-C Implementation Details
The training is done in two stages. In the first stage, we fit the parameters of encoders and decoders by maximizing the VAE ELBO in Eq. 2. In the second stage, we fit the parameters of the score model to minimize DDM loss in Eq. 6. While it is possible to train all the networks in a single stage, it is easier and faster to train the diffusion model once the latent space of the VAE is frozen. In single-stage training, the constantly changing posterior distribution at the encoder means that the optimization of the weights of the de-noising score network is wasteful until VAE training has converged sufficiently. To train our networks, we use objects and grasp annotations from the ACRONYM [17] dataset. We use train-test splits defined in [12]. We empirically choose the ELBO modifying hyperparameter $\lambda$, which is annealed linearly from $71e-$ 7 to $0.1$ for $50\%$ of the training steps and then held constant until the end of training. This ensures that the KL term does not vanish in the early stages of the training while also enforcing lower regularization to prioritize grasp pose reconstruction during the later stages.
The point cloud encoder is based on Point-Voxel CNN (PVCNN) architecture [18] which is more efficient than conventionally used PointNet architecture [18]. The input to the point cloud encoder is fixed at 1024 points, to balance memory usage with point sampling density. The output of the point cloud encoder is a latent ($\mathbf{z}_{pc}$) of size 128. The grasp pose encoder and decoder are 1D convolutional neural networks with residual blocks. Conditioning is done using FiLM [19], which uniquely scales and shifts the feature maps at every residual block based on the conditioning features. The decoder network is the same as the encoder but reversed, whose output is the reconstruction ($\mathbf{x}^{*}_{h}$) of the input. The noise prediction network for diffusion is a residual network similar to the grasp encoder and is conditioned on both the diffusion time step ($t$) and the shape latent ($\mathbf{z}_{pc}$).
@@
We use the representation of the grasp pose ($\mathbf{h}$) given in Eq. 9, where $\mathbf{t}$ is the translation vector while $\mathbf{a}$ is the Modified Rodrigues Parameters (MRP[20] for orientation. Eq. 9 also relates MRP with quaternions where $q_{w}$ is the quaternion scalar component. In contrast to regressing pose matrices or quaternion vectors that have one or more dependent parameters, the three parameters of MRP s can be regressed independently. We encourage readers to check the details of the implementation in our code release.

### v2

@@ Abstract
Vision-based grasping of unknown objects in unstructured environments is a key challenge for autonomous robotic manipulation. A practical grasp synthesis system is required to generate a diverse set of 6-DoF grasps from which a task-relevant grasp can be executed. Although generative models are suitable for learning such complex data distributions, existing models have limitations in grasp quality, long training times, and a lack of flexibility for task-specific generation. In this work, we present GraspLDM- a modular generative framework for 6-DoF grasp synthesis that uses diffusion models as priors in the latent space of a VAE. GraspLDM learns a generative model of object-centric $SE(3)$ grasp poses conditioned on point clouds. GraspLDM’s architecture enables us to train task-specific models efficiently by only re-training a small denoising network in the low-dimensional latent space, as opposed to existing models that need expensive re-training. Our framework provides robust and scalable models on both full and single-view point clouds. GraspLDM models trained with simulation data transfer well to the real world without any further fine-tuning. Our models provide an 80% success rate for 80 grasp attempts of diverse test objects across two real-world robotic setups.
Index Terms:
generative modeling, robotic grasping, grasp synthesis, diffusion models
[…]
=-21pt
@@ I Introduction
Robotic grasping of unknown objects from visual observations is an integral skill for autonomous manipulation systems in the real world. Grasping is the fundamental manipulation task of generating restraining contacts on an object. An essential part of this task is grasp synthesis, which involves reasoning about a set of grasp configurations that are possible around an object. From this set of possible grasps, a subjectively good grasp can then be chosen based on grasp quality, task context, and kinematic feasibility. Real-world grasping requires a robot to determine grasping location and configuration from incomplete perceptual observations. Cameras are convenient as they are cost-effective and provide dense information in their field of view. However, the pixel-level information from a camera is unstructured and cannot be trivially mapped to geometric information that is helpful for manipulation, especially in unknown environments with object models that are not known apriori.
Figure 1: GraspLDM models trained on synthetic data successfully transfer to the real world and provide stable 6-DoF grasps from single-view RGB-D data in the presence of workspace and motion planning constraints.
For general-purpose robotic grasping, providing a single grasp configuration corresponding to an observation is not a suitable approach, as the grasp configuration might be kinematically unfeasible and contextually unsuitable for a task. On the other hand, we must develop systems that are task-agnostic but easy to adapt for specific tasks. This can be accomplished by designing vision-based grasp synthesis systems that more generally reason about a set of grasps around an object, further allowing selection based on a metric. However, vision-based grasp synthesis is a challenging problem for three main reasons. First, there are infinite possible grasps on any object and the distribution of good grasps can be complex, with multiple modes and discontinuities. Second, the grasp generation process has to rely on unstructured visual observations from the sensors. Finally, it needs to generalize to an arbitrary set of unknown objects that a robot may encounter in the real world. To alleviate some of these issues, contemporary works have increasingly relied on data-driven approaches [1]. This is in contrast to analytical methods for grasp synthesis [2], which assume the knowledge of an object’s physical and geometric models and subsequently execute high-dimensional search satisfying an optimality criterion. Apart from the computational complexity, the grasps generated through such analytical grasp quality optimization could also transfer poorly to the real world [3].
Instead of the dynamics of hand-object contacts, data-driven grasp synthesis approaches focus on perceptual processing and generalizable representations [1]. Early works focused on transferring grasps by using some notion of familiarity to a known reference [4, 5]. Subsequently, machine learning drove significant progress as the mapping from visual data to grasp configurations could be modeled using learned abstract representations. The Cornell dataset and the grasp rectangle representation introduced in [6] stemmed a line of research [7, 8] significantly advancing the state-of-the-art in grasping unknown objects. However, these approaches restrict the Degrees of Freedom (DoF)of grasping to a plane perpendicular to the image plane, typically used for top-down grasping from a table-top. In contrast, synthesizing 6-DoF grasp configurations on unknown objects remains a challenge [9].
Learning-based 6-DoF grasp synthesis can be accomplished by supervised learning on annotated grasp data, reinforcement learning in relevant environments, or using exemplar learning by imitation. Reinforcement learning methods suffer from sample complexity and hyperparameter sensitivity [10], while imitation learning is bottlenecked by human demonstration quality. In comparison, supervised learning has shown increased performance and generalization, with the availability of large and high-quality open-source datasets with dense grasp annotations [11, 12]. Within supervised learning, however, most approaches are based on direct regression of grasp poses or grasp quality. Fewer works [13, 14] employ generative modeling. Generative modeling approaches learn the structure of the input data and efficiently model the data-generating distribution. Generative models [15, 16, 17] have consistently improved performance in real-world generation and synthesis across domains like natural language and computer vision. In particular, a new class of generative models called Denoising Diffusion Model (DDM)have recently shown state-of-the-art performance on likelihood-based metrics and output sample quality [18, 16, 19]. Given the successful application of DDM s to complex generation tasks for images [16] and point clouds [20], we investigate their potential for improving generative modeling in robotics.
@@
We focus on the problem of generating object-centric 6-DoF grasps using point cloud observations. We assume a parallel jaw gripper and parameterize a grasp by its $SE(3)$ pose. The objective is to learn a representation of object-centric 6-DoF grasps that can generalize to unknown objects. Learning this representation is difficult, as the model must reason about an arbitrary set of grasp poses for each object in the $SE(3)$ space. Deep generative models are suitable for this problem, as they are designed to learn complex data-generating distributions efficiently. 6DoF-Graspnet [13], an early generative grasp synthesis model, learns the distribution of object-centric grasps in a continuous latent space of a conditional Variational Autoencoder (VAE). However, the VAE exhibits poor sampling quality and requires an additional iterative post-processing stage to improve grasp samples. The shortfalls of the VAE to learn effective representations could be attributed to the prior gap problem in VAE[21]. The prior gap problem refers to the gap between the posterior distribution modeled by the VAE encoder and the assumed prior, which results in poor quality samples when decoding random latents from the prior. On the other hand, diffusion models [17, 16] can overcome these problems and are simpler to train. In particular, the flexible generation architecture of latent diffusion [22] is especially appealing.
Figure 2: GraspLDM uses a denoising diffusion model in the latent space of a VAE to improve grasp generation performance. It also enables injection of task-conditional guidance in a modular manner.
@@
In this work, we propose GraspLDM - a learning-based generative modeling approach for 6-DoF grasp synthesis. GraspLDM learns the distribution of successful grasps on object point clouds using latent diffusion that transfers to the real world (Fig. 1). In GraspLDM, a diffusion model is trained efficiently inside the low-dimensional latent space of a VAE. This diffusion model acts as a prior and bridges the gap between the VAE prior and the posterior distribution fit during the training. Using diffusion in the latent space improves the quality of sampled grasps and avoids the representational complexity of operating directly on pairs of point clouds and grasp pose. It especially enables two aspects of model flexibility that are not available in existing generative models for grasp synthesis using VAE [13] and diffusion [23] alone. First, task-specific conditional guidance can be provided solely in the latent space without re-training the VAE encoder and decoder. Second, multiple diffusion models can be trained efficiently and plugged inside this low-dimensional latent space. Further, the choice of latent diffusion structure is justified by the high dimensionality of representation required to directly operate on a pair of point cloud and grasp in a vanilla diffusion model.
@@ II Related Work
Grasp synthesis was initially studied using analytical methods [2], which rely on the knowledge of object models and properties that are rarely available in the real world. The complexity of the search problem and the dependence on models paved the way for methods that learn directly from data. To learn from data, the focus changed from contact modeling to designing an effective visual processing system and choosing robust representations. Such works initially relied on primitive decomposition [5], demonstrations [24], template matching [25] or transferring some notion of grasp familiarity [4]. More recently, grasp synthesis has been framed as the problem of finding stable grasp poses directly from sensor data, subject to a gripper. Data-driven and learning-based methods have made notable progress, especially on 4-DoF planar grasp synthesis using parallel-jaw and vacuum grippers. In contrast, generalizable 6-DoF grasp synthesis is an active research topic [9] with two broad classes: discriminative and generative approaches. Discriminative approaches rely on a manual pose sampling strategy and learn to discriminate good grasps from bad grasps using a quality measure [7]. Generative approaches instead directly learn to generate grasp poses from an observation [26, 13] with implicit sampling. Since grasps can be sampled directly from the learned model [13] these tend to be efficient. Some works on 6-DoF grasping also address auxiliary tasks to improve grasp generation, like shape completion and 3D reconstruction [27]. In this work, we present a generative approach that learns a continuous distribution of object-centric grasps conditioned on point clouds in an end-to-end manner.
@@ II-C Denoising Diffusion Models
GraspLDM utilizes a diffusion model based on Denoising Diffusion Probabilistic Models (DDPM)[16]. Score-based Generative Model (SGM[17] and Stochastic Differential Equations (SDE)are equivalent formulations of the diffusion process with notable implementation differences. For grasp pose generation, [23] uses the SGM formulation and learns a scalar field representing the denoising score. While their models effectively refine randomly sampled SE(3) grasp poses to low-cost (good grasp) regions, the analysis only presents learning on full point clouds of a single category (Mugs). Real-world tests on Mugs in [23] use the ground truth pose of the object. Further, Langevin-type sampling is slow and the model needs to be retrained for downstream tasks. In contrast, GraspLDM is a more flexible formulation to [23] that operates a diffusion model on a vector field in the continuous latent space of a VAE. Our analyses and validation are also more extensive using full and partial point clouds from a large object category set, complemented with real-world transfer tests. GraspLDM can also use fast samplers like Denoising Diffusion Implicit Model (DDIM)as a drop-in replacement post-hoc with minimal performance loss. Further, GraspLDM can be extended to gripper parameterizations beyond the pose, in contrast to [23]. Diffusion models have also been utilized for dexterous grasp synthesis [28], but rely on the availability of the object model.
@@ III Grasp Latent Diffusion Models
Since we do not have access to the real latent space, computing the likelihood of data over the whole latent space in Eq. 1 is intractable. To overcome this, we use a VAE formulation for learning our generative model. VAE’s encoder compresses the point cloud-grasp pair input to a continuous latent space with a prior distribution. However, the approximate posterior of the encoder does not match the prior perfectly, especially if it prioritizes data reconstruction for grasp accuracy. To tackle this, we use a diffusion model in its latent space. Together, they form the latent diffusion model for grasp generation that we call GraspLDM. The diffusion model simultaneously improves sample quality by bridging the prior gap in the latent space and enables the introduction of latent space guidance with task context. The following sections describe the GraspLDM design in detail.
@@ III-A Conditional Variational Autoencoder
In Eq. 2, the first term is the likelihood of the decoder reconstructing the input data. The second term measures the divergence between the approximate posterior distribution $q_{\psi}(\mathbf{z}_{h}|H,\mathbf{x}_{pc})$ of the encoder and the assumed prior, $\mathbf{z}\sim\mathcal{N}(0,1)$. $\lambda$ is the weighting hyperparameter that controls the strength of KL regularization in the latent space. A constant $\lambda=1$ provides the strict ELBO objective. However, this frequently results in the KL divergence term ($D_{KL}$) becoming vanishingly small early in the training. When this happens, the conditional decoder becomes auto-regressive and the latent has no effect on the output [21]. To avoid this, we use linear annealing on the $\lambda$ parameter.
@@ III-B Latent Diffusion
In principle, sampling from VAE’s prior and decoding it should be sufficient to generate successful grasps. However, decoding a latent drawn from the prior often results in lower sample quality because of the prior gap problem introduced in Section I. To alleviate this issue, we propose to use a Denoising Diffusion Model (DDM)in the latent space. We justify the choice of latent diffusion structure from two observations. First, a simple DDPM would need to encode a point cloud at each step and use a high-dimensional intermediate representation, increasing computation for each denoising step. Alternatively, separating the point cloud encoder from the diffusion requires separate expensive training of a point cloud auto-encoder. Further, such a model needs to be re-trained for newer task conditionings.
A DDM consists of two processes: a forward diffusion process and a reverse denoising process. In this work, we use the standard discrete-time DDPM [16], where the forward process is a linear Markov chain of deterministic Gaussian kernels in Eq. 3a. The noise added at each time-step $t\in[1,T]$ is pre-defined by a variance schedule $\beta_{t}$. We use linearly increasing $\beta_{t}$ such that the distribution $q(\mathbf{z}_{T})$ converges to the standard normal distribution at the final forward time step $T$. In the forward diffusion process, a noisy sample at any time step can then be obtained by Eq. 3b.
@@
Intuitively, the forward diffusion process adds noise to input data until the data distribution transitions to a simple prior distribution. Consequently, recovering the data distribution from the prior requires a time reversal of the forward process from $t=T$ to $t=0$. This reverse diffusion process uses a learned score network ($\theta$) which de-noises a sample, in our case the latent $\mathbf{z}_{h}$, from $t$ to $t-1$. For notational simplicity, we use $\mathbf{z}_{t}$ to refer to conditional grasp latent $\mathbf{z}_{h}$ at time $t$. Using the reverse diffusion kernel in Eq. 4, the distribution of the de-noised sample at time-step $t-1$ can be modeled by the mean $\mu_{\theta}(z_{t},t)$ and a known variance $\sigma_{t}^{2}$. More simply, we can re-parameterize the mean and instead learn a network to directly predict the noise $\bm{\epsilon}_{\theta}(\mathbf{z}_{t},t)$ to be removed at each step [16]. The denoising score model $\bm{\epsilon}_{\theta}$ predicts the noise to be removed, conditioned on the current time-step, while the parameter weights ($\theta$) are shared across all the time steps. Then, the sampling can be done using Eq. 5, where $\sigma_{t}$ can be simplified to use variance similar to the forward process i.e. $\beta_{t}$. DDMs with this formulation can be trained to maximize an ELBO-like objective. Given the simplified formulation in [16] and adapting it to our latent diffusion architecture, this objective $\mathcal{L}_{D}(\theta)$ can be written as Eq. 6.
@@ III-C Implementation Details
The training is done in two stages. In the first stage, we fit the parameters of encoders and decoders by maximizing the VAE ELBO in Eq. 2. In the second stage, we fit the parameters of the score model to minimize DDM loss in Eq. 6. While it is possible to train all the networks in a single stage, it is easier and faster to train the diffusion model once the latent space of the VAE is frozen. In single-stage training, the constantly changing posterior distribution at the encoder means that the optimization of the weights of the denoising score network is wasteful until VAE training has converged sufficiently. To train our networks, we use objects and grasp annotations from the ACRONYM [11] dataset. We use train-test splits defined in [26]. We empirically choose the ELBO modifying hyperparameter $\lambda$, which is annealed linearly from $71e-$ 7 to $0.1$ for $50\%$ of the training steps and then held constant until the end of training. This ensures that the KL term does not vanish in the early stages of the training while also enforcing lower regularization to prioritize grasp pose reconstruction during the later stages.
The point cloud encoder is based on Point-Voxel CNN (PVCNN) architecture [29] which is more efficient than conventionally used PointNet architecture [29]. The input to the point cloud encoder is fixed at 1024 points, to balance memory usage with point sampling density. The output of the point cloud encoder is a latent ($\mathbf{z}_{pc}$) of size 128. The grasp pose encoder and decoder are 1D convolutional neural networks with residual blocks. Conditioning is done using FiLM [30], which uniquely scales and shifts the feature maps at every residual block based on the conditioning features. The decoder network is the same as the encoder but reversed, whose output is the reconstruction ($\mathbf{x}^{*}_{h}$) of the input. The noise prediction network for diffusion is a residual network similar to the grasp encoder and is conditioned on both the diffusion time step ($t$) and the shape latent ($\mathbf{z}_{pc}$). For this purpose, we first embed the time step scalar to the desired dimensions using sinusoidal positional embedding. The shape latent is then projected onto the same dimension and added to the time embedding, before applying it to the features at every block.
@@
We use the representation of the grasp pose ($\mathbf{h}$) given in Eq. 7, where $\mathbf{t}$ is the translation vector while $\mathbf{a}$ is the Modified Rodrigues Parameters (MRP[31] for orientation. Eq. 7 also relates MRP with quaternions where $q_{w}$ is the quaternion scalar component. In contrast to regressing pose matrices or quaternion vectors that have one or more dependent parameters, the three parameters of MRP s can be regressed independently.
We train our models on both full and partial point clouds. In each case, the point clouds are augmented online with random rotations, point jitter, and point dropout. Random rotation is sampled by selecting a random unit vector (axis) and a random magnitude (angle). Point jitter randomly perturbs the 3d coordinate of the points with a standard deviation of 1cm in all directions around the original to emulate jitter artefacts of a noisy sensor stream. Point dropout randomly reduces the availability of unique points by removing up to 40% of the input points. Since the input point cloud has to be of a fixed shape, the removed points are replaced by duplicating existing points. To avoid relying on an explicit canonical frame, the point cloud and grasp poses are both expressed in a frame with the origin at the centroid of the input point cloud.
For the first term of the ELBO (Eq. 2) loss that measures reconstruction, we use L2 loss between inputs and outputs of the 6-parameter grasp pose. For learning rate decay, we scale down the learning by 0.1 every 1/3rd of the total training steps, going from 1e-3 to 1e-5 to avoid large optimization steps in later stages of the training. For the diffusion model, the noise variance ($\beta_{t}$) for forward diffusion is set to a linear schedule $\beta$ with $\beta_{0}=5e-5$ and $\beta_{T}=1e-3$. The denoising diffusion formulation closely follows the original DDPM formulation with 1000 time steps in the nominal configuration.